# Emergent spin order and steady-state superradiance in one-dimensional baths

### v1

@@ Appendix B STEADY-STATE ANSATZ FOR THE WAVEGUIDE
$+P_{nm}(R,L)e^{-{\rm i}k(z_{n}+z_{m})}\int_{0}^{2\pi}\frac{d\varphi_{R}}{2\pi}\int_{0}^{2\pi}\frac{d\varphi_{L}}{2\pi}e^{{\rm i}(\varphi_{R}-\varphi_{L})}$
$+P_{nm}(L,R)e^{{\rm i}k(z_{n}+z_{m})}\int_{0}^{2\pi}\frac{d\varphi_{R}}{2\pi}\int_{0}^{2\pi}\frac{d\varphi_{L}}{2\pi}e^{{\rm i}(\varphi_{L}-\varphi_{R})}.$
@@
Appendix C FURTHER DETAILS ON LINEWIDTH
Here, we further elaborate on the calculation of the minimum achievable linewidth and its scaling with atom number $N$, and outline the main constraints of current techniques.
In the absence of symmetries, the numerical complexity for computing the linewidth of any field using the TWA scales as $\mathcal{O}(MN^{3})$ [Guardiola25], where $M$ is the number of trajectories required for convergence. The procedure requires the calculation of the multi-time field correlator and fitting it to an exponential $e^{-\tau\Delta\nu}$. Numerical exploration of this fitting indicates good agreement within the expected limitations of the TWA, as the correlator does not vanish for $\tau\Gamma_{\rm 1D}\gg 1$ due to the inherent noise of the approximation.
These calculations are performed for several values of $w$ to estimate the minimum linewidth. To reduce computational cost and mitigate the effects of outlier and noisy realizations, we adopt the following strategy: instead of performing a dense scan, we compute the linewidth for only 10-15 selected pumping strengths (not necessarily uniformly spaced nor including the optimum), and then fit to a second-order polynomial. Alternative models, including higher-order polynomials or other non-linear functions, could be used to estimate the minimum, although resulting in similar values.

### v2

@@ Appendix B STEADY-STATE ANSATZ FOR THE WAVEGUIDE
$+P_{nm}(R,L)e^{-{\rm i}k(z_{n}+z_{m})}\int_{0}^{2\pi}\frac{d\psi_{R}}{2\pi}\int_{0}^{2\pi}\frac{d\psi_{L}}{2\pi}e^{{\rm i}(\psi_{R}-\psi_{L})}$
$+P_{nm}(L,R)e^{{\rm i}k(z_{n}+z_{m})}\int_{0}^{2\pi}\frac{d\psi_{R}}{2\pi}\int_{0}^{2\pi}\frac{d\psi_{L}}{2\pi}e^{{\rm i}(\psi_{L}-\psi_{R})}.$
@@
Appendix C LINEWIDTH ANALYSIS
Here, we provide further details on the calculation of the linewidth, the scaling of the minimum linewidth with the atom number $N$, and the estimation of error bars. We also discuss the main limitations of the numerical techniques employed.
In the absence of symmetries, the numerical complexity for computing the linewidth of any field using the TWA scales as $\mathcal{O}(MN^{3})$ [26], where $M$ is the number of trajectories required for convergence. To estimate the linewidths shown in Fig. 6(a) of the main text, we compute the two-time correlator of either the cavity field or the right-propagating field and fit it to an exponential decay of the form
$\langle\hat{E}^{-}(t_{ss}+\tau)\hat{E}^{+}(t_{ss})\rangle\sim e^{-\tau\Delta\nu/2}$ (47)
Although fluctuations inherent to the TWA induce a noise floor at late times, preventing the correlator from vanishing completely as $\tau\rightarrow\infty$, we find that the fitting procedure reliably captures the initial exponential decay of the correlator. The error bars in Fig. 6(a) arise from the stochastic nature of the TWA simulations and correspond to three standard deviations estimated from the fitting procedure.
To determine the minimum linewidth, these calculations are repeated for several values of the incoherent pumping rate $w$. To reduce computational cost and mitigate the impact of noisy or outlier realizations, we do not perform a dense scan in $w$. Instead, we evaluate the linewidth at 10 to 20 selected pumping strengths, which are not necessarily uniformly spaced and may not include the value of $w$ that yields the absolute minimum linewidth. To estimate the minimum from these discrete samples, we fit the data to a LogSumExp function, $\frac{1}{p_{0}}\log\left(\exp{(p_{0}(p_{1}w+p_{2}))}+\exp{(p_{0}(p_{3}w+p_{4}))}\right)$, from which we obtain the minimum linewidth. Alternative fitting models, including polynomials and other non-linear functions, can be used, resulting in similar estimates for the minima. The error bars in Fig. 6(b) likewise represent three standard deviations.